# Machine Learning with Physics Knowledge for Prediction: A Survey

### v1

@@ 3.5 Latent Variable Models for Dynamical Systems
Despite its theoretical concepts, the practical implementation of Koopman theory is ongoing (Brunton et al., 2021). While promising, autoencoder methods typically serve as dimensionality reduction techniques, conflicting with the concept of a potentially infinite-dimensional latent space. Nevertheless, they have shown remarkable performance in finding low-dimensional embedding for tasks involving high-dimensional solution spaces (Brunton et al., 2021). Furthermore, several works including Lusch et al. (2018), have rendered the linear latent dynamics as state-dependent resulting in locally-linear non-linear dynamics. Although these architectures perform well on high-dimensional nonlinear fluid flow, they loosen the Koopman invariance assumption due to the operator’s dependence on the latent state. These results motivate the discussion of architectures that employ dimensionality reduction while relaxing the linearity assumption for the dynamics in latent space.
Nonlinear latent dynamics. Modeling a dynamical system without assuming latent linear dynamics removes the inductive bias but yields a more expressive parametric model. In the context of this survey, these models are less relevant as their only inductive bias is the Markovian structure of the data, but these models perform well on a range of complex tasks, such as modeling natural language (Sutskever et al., 2014), music (Roberts et al., 2018) and video data (Babaeizadeh et al., 2018). One relevant line of work investigated how to implement these models with physically-structured latent space, like the models discussed in section 3.3. These include position-velocity encoders (Jonschkowski et al., 2017), Hamiltonian generative networks (Toth et al., 2020) and Newtonian variational autoencoder (Jaques et al., 2021). More recently, significant progress has been made in machine learning for weather prediction using deep latent variable models. While these models are primarily data-driven and do not have physically-interpretable latent spaces, their performance has been achieved using several small but crucial physics-informed design decisions that we will discuss here.
@@
Weather forecasting traditionally relies on computationally intensive physically-derived models, often combined in ensembles, to mitigate uncertainties arising from chaotic atmospheric dynamics and measurement limitations. Despite advancements, the accuracy of these methods for medium-range, regional, and global climate modeling is dictated by the computational resources and model fidelity. Additionally, they cannot leverage extensive historical weather data beyond optimizing model parameters with system identification. Moreover, the complexity of weather dynamics requires combining models for each component (e.g. clouds, pollution, etc), and jointly tuning these models to improve net prediction accuracy is challenging/ Machine learning approaches offer promise by directly predicting measurements from historical datasets, circumventing some computational challenges. Recent years have witnessed a significant advancement in machine learning methods for weather prediction, enabled by large capacity models, hardware accelerators, and clean historical datasets (Voosen, 2023). We discuss three prominent nonlinear latent variable models in this context.
@@
Bi et al. (2023) proposed Pangu-Weather. Like FourCastNet, this model builds off vision transformers, but instead of using AFNOs, proposes a novel three-dimensional (3D) Earth-specific transformer architecture (3DEST) to better model weather correlation with latitude, longitude, and altitude. The key design decision is in the positional encoding used by the 3DEST, which encodes absolute position coordinates rather than relative ones and incorporates translation invariance. This encoding uses 527 times more embedding parameters, but the authors did not observe issues with prediction or optimization. The model also employs ‘hierarchical temporal aggregation’ to address error accumulation in long-horizon forecasts. To mitigate this source of error, the authors learn models that are trained on 1-, 3-, 6- and 24-hour intervals so that long-horizon forecasts can be computed using an aggregation of models, using 3-4 predictions at most. However, with each 3D ViT having 64M parameters, this aggregation strategy results in a 256M parameter model, which presents engineering issues at deployment.
@@
Lam et al. (2023) introduced GraphCast, a graph neural network (GNN) approach for medium-range forecasting. GraphCast utilizes a hierarchical graph structure over the globe, resembling 3DEST, with six levels of resolution. The model’s encoder projects Earth-oriented grid measurements onto this multi-mesh representation while the decoder aggregates nearby mesh values back to the grid. Operating on a 40,962-node multi-mesh, the GNN conducts simulation via simultaneous message passing at various spatial resolutions. Despite making 6-hour predictions, during training, it is evaluated autoregressively for up to 3 days. Remarkably, GraphCast can predict a 10-day forecast in under a minute at 0.25 global resolution with approximately 40 million parameters. Comparative evaluations suggest GraphCast outperforms Pangu-Weather across most measurement modalities, possibly due to the GNN’s inductive bias, which resembles finite-element PDE solvers and propagates computations coherently in space. This behavior contrasts with ViT architecture, which needs latent self-attention mechanisms to learn such computations. However, the graph architecture does complicate some aspects of forecasting, such as downscaling and regional prediction.
@@
Despite progress, several challenges persist in current weather forecasting models. A significant hurdle is their reliance on ‘reanalysis’ data derived from complex Bayesian state estimation processes, which combines meteorological measurements with simulation models. This reliance on processed data prevents these ML models from forecasting in real time using raw data as desired. Additionally, challenges include scaling models to go beyond the 0.25 training data and match the high spatial resolution of 0.1 of top NWP systems, extending forecasting horizons, and increasing model capacity. There is also the question of the extent to which physics-based elements from NWP can be combined with ML models to enhance their performance, given that NWP forecasts are inherently physically plausible, while current ML-based forecasts lack this feature. This question is an active topic of study, e.g., (Kochkov et al., 2024). The next section explores examples of incorporating physics-informed elements into model learning.
3.6 Bridging System Identification and Function Approximation
@@
System identification approaches lie on a spectrum depending on modeling assumptions. If a lot of domain knowledge is assumed, the dynamics model will be highly structured. If correct, it will generalize well but will most likely be inflexible to un-modelled phenomena. On the other hand, one could model with no modeling assumptions and use a black-box function approximator [^3] One could argue that using many popular function approximators, such as Gaussian processes, come with a smoothness assumption. These models have no generalization guarantees and, therefore, require expansive datasets and regularization to ensure good modeling accuracy across the domain of interest. To achieve an effective combination of these two approaches, one would need to balance structural biases that facilitate generalization with adaptivity to the data that prevents the biases from leading to underfitting (Ljung, 2010; Lutter et al., 2021). This section reviews these approaches, which build on models discussed earlier.
@@
Physics-informed neural networks can incorporate an additional dataset as an additional loss term in the objective. In addition to residual and boundary loss terms, the combined objective (equation  (13)) has an additional dataset loss similar to the boundary loss term (Raissi et al., 2017a; b). Moreover, since the PINNs objective is quite general to any differentiable function approximator, Wang et al. (2021c) extend DeepONets by adding the PINNs residual loss to the data-driver operator loss as a way of encoding physics knowledge.
@@
3.7 Invariance-Driven Priors
Physical models in the form of ODEs, PDEs, or algebraic equations are often invariant to specific types of operations or transformations. For instance, many dynamical processes such as molecular, rigid body, elasto or fluid dynamics, chemical reaction kinetics, etc. are modelled under the assumption of energy conservation and the laws of thermodynamics (Cueto and Chinesta, 2023). This also translates to continuum material models, which should be invariant to the orientation of the coordinate system and consider the orientation of the material’s microstructure. Likewise, in computer vision, object detection should be invariant to the position, orientation, and scale of the object and graph-based representations, e.g., in the context of molecular interactions or particle systems (Sanchez-Gonzalez et al., 2020), should often also be invariant to rotations, translations, reflections, and permutations. To discover invariants and conservation quantities from datasets using machine learning, Wetzel et al. (2020) proposed a “Siamese twin” neural network architecture, which aims to classify data points as related to the same invariant value. Similarly, Ha and Jeong (2021) introduced an intuitive noise-variance loss function for training a NN to classify data points related and non-related to the same invariant value.
Mathematically defining and incorporating relevant invariance and equivariances into machine learning models has led to several advancements in applying machine learning to complex physical domains, e.g., (Cohen and Welling, 2016; Fuchs et al., 2020; Satorras et al., 2021; Hoogeboom et al., 2022; Weiler et al., 2023). Denoting a set of transforms as $\mathsf{T}$, the $\mathsf{T}$-invariance of a function or operator can be expressed as:
@@ 3.7 Invariance-Driven Priors
(model-based) $f_{\bm{\theta}}({\bm{x}})$ $=f_{\bm{\theta}}(T({\bm{x}}))\quad,\forall T\in\mathsf{T},$ (42)
(feature-based) $f_{\bm{\theta}}(\phi({\bm{x}}))$ $=f_{\bm{\theta}}(\phi(T({\bm{x}})))\quad,\forall T\in\mathsf{T},$ (43)
@@
In summary, material modeling is subject to several types of invariance requirements such as thermodynamic consistency, rotational invariances, and functional requirements. It has already been demonstrated through various techniques that this prior knowledge can be successfully incorporated into ML frameworks for constitutive modeling. In particular, ML models can be fully physics-augmented, i.e., the requirements can be embedded into the model architectures by careful choice of features, outputs, and NN structure. This generally greatly improved the model performance, generalization, robustness, and interpretability and reduced data requirements.
@@
Symmetries and energy conservation are also important aspects of the modeling of dynamic systems. As already discussed in section 3.3, the equations of motion of a dynamical system can be formulated in an energy-conserving, i.e., thermodynamically consistent, fashion by deriving them through the Lagrangian or Hamiltonian formalisms (Lutter and Peters, 2023; Cranmer et al., 2020b; Greydanus et al., 2019). In these approaches, the invariance is model-based and incorporated through the model conception as the (Lagrangian ${\mathcal{L}}$, Hamiltonian ${\mathcal{H}}$, …) potentials are learned/parameterized by a NN – and not the dynamics themselves in the form of an ODE, e.g., through a neural ODE, see section 3.1. This concept has also been applied to Koopman operators, c.f. section 3.5, for the data-based discovery of conservation laws in Kaiser et al. (2018) and could be further extended to neural operators.
However, energy conversation is usually an idealization of real system behavior, and dissipative effects are present in most real-world systems. Dissipation behavior can be considered in a similar fashion through application of formalisms such as the port-Hamiltonian (Desai et al., 2021; Neary and Topcu, 2023), see equation  (20), pseudo-Hamiltonian (Eidnes et al., 2023), or GENERIC (General Equation for the Non-Equilibrium Reversible–Irreversible Coupling) frameworks (González et al., 2019; Hernández et al., 2021), which is structurally very similar to the GSM framework mentioned above for material modeling.
@@ 4.1 Multi-task Learning
Thanasutives et al. (2021) propose an MTL approach of solving the original PDE besides to an auxiliary related PDE with different coefficients. In this work, they integrate MTL at the architecture and optimization levels. For the architecture, cross-stitch (Misra et al., 2016) is employed to share information between two parallel baseline networks for original and auxiliary tasks. At each layer, a cross-stitch unit is used to linearly combine the input activations from the two networks to allow flexible sharing of representations between the two branches. The authors evaluate two MTL techniques for learning, an uncertainty-weighting loss (Kendall et al., 2017) and gradients surgery (Yu et al., 2020). The uncertainty-weighted physics-informed loss function (Uncert) is defined as a combination of task-specific PINNs loss functions weighted by the uncertainty of each PDE solution. The loss function is defined as,
@@
where $N$ refers to the number of tasks, $\sigma_{i}$ is a gradient-based trainable parameter that indicates the uncertainty of each PDE solution, and $\mathcal{L}_{\textrm{PINN}}$ is a similar loss function as in equation  (13). ‘Gradient surgery’ is implemented using PCGrad (Yu et al., 2020) to modify the average of the unweighted PINNs loss functions for mitigating conflicting gradients between tasks. PCGrad projects the loss gradient of task $i$ onto the normal vector of task $j$’s loss gradient and vice versa. The model parameters are updated using $\delta_{\textrm{PC}}\theta$ shown below,
@@
where ${\bm{g}}^{i}_{\textrm{PC}}$ is the modified gradient of task $i$ after applying PCGrad. The proposed algorithm provides a better generalization on high non-linearity domains in terms of lower error rates on unseen data points in comparison to related baselines.
Zou and Karniadakis (2023) propose a new PINN architecture that suites MTL. In this work, a multi-head architecture is presented, named multi-head physics-informed neural network (MH-PINN), where a body $\Phi$ is shared by a collection of tasks (e.g., ODEs/ PDEs) while task-specific heads $H_{k},k=1,...,N$ are learned separately for each model. The body consists of non-linear hidden layers, while linear layers are used for the different heads. Consequently, the body provides a set of basis functions for each head to generate task-specific solutions. The body $\Phi:\mathbb{R}^{d_{{\bm{x}}}}\rightarrow\mathbb{R}^{d}$ is a function parameterized with parameter $\theta$, and each head $H_{k}$ is a linear function. The task-specific solution is defined as $\hat{{\bm{u}}}_{k}({\bm{x}})=H_{k}\circ\Phi(x),\forall x\in\Omega$. Given an MTL dataset $\{\mathcal{D}_{k}\}_{k=1}^{N}$, the MTL loss function is defined as follows:
@@
Similarly, Wang et al. (2023) ease the training of PINNs by employing a multi-task optimization paradigm where multiple auxiliary tasks are solved together with a main task. Knowledge transfer from one task facilitates solving others. The problem of automatic traffic modeling is considered, which has multiple modalities such as traffic speed and density. Given several datasets measuring different traffic scenarios and MTL over datasets and modalities, a similar PINN architecture, as in Zou and Karniadakis (2023), is trained with a shared body network and only a final layer per task. This significantly increases the data available to train the initial PINN layers, which build an internal feature representation. The MTL ensures that this feature representation is useful across traffic conditions and modalities. As a drawback, the model requires an adaptive training scheme in order for the multi-task aspect to be effective, which significantly increases the implementation complexity.

### v2

@@ 3.5 Latent Variable Models for Dynamical Systems
Despite its theoretical concepts, the practical implementation of Koopman theory is ongoing (Brunton et al., 2021). While promising, autoencoder methods typically serve as dimensionality reduction techniques, conflicting with the concept of a potentially infinite-dimensional latent space. Nevertheless, they have shown remarkable performance in finding low-dimensional embedding for tasks involving high-dimensional solution spaces (Brunton et al., 2021). Furthermore, several works , including Lusch et al. (2018), have rendered the linear latent dynamics as state-dependent resulting in locally-linear dynamics. Although these architectures perform well on high-dimensional nonlinear fluid flow, they loosen the Koopman invariance assumption due to the operator’s dependence on the latent state. These results motivate the discussion of architectures that employ dimensionality reduction while relaxing the linearity assumption for the dynamics in latent space , i.e., nonlinear latent dynamics .
Nonlinear latent dynamics. Modeling a dynamical system without assuming latent linear dynamics removes the inductive bias but yields a more expressive parametric model. In the context of this survey, these models are less relevant as their only inductive bias is the Markovian structure of the data, but these models perform well on a range of complex tasks, such as modeling natural language (Sutskever et al., 2014), music (Roberts et al., 2018) and video data (Babaeizadeh et al., 2018). One relevant line of work investigated how to implement these models with physically-structured latent space, like the models discussed in section 3.2. These include position-velocity encoders (Jonschkowski et al., 2017), Hamiltonian generative networks (Toth et al., 2020) and Newtonian variational autoencoder (Jaques et al., 2021). Another noteworthy line of research, discussed in section 3.4, concentrates on operator learning. DeepONets employ an autoencoder architecture to model nonlinear operators within a finite-dimensional latent space (Lu et al., 2019).
Encoder-decoder models can also be combined with neural ODEs (section 3.1). Dynamics-aware implicit neural representations (DINo, Yin et al. (2023)) uses an implicit neural representation (INR) as a decoder. INRs describe parametric functions trained to describe a domain, such as those already seen in PINNs (section 3.3) and also neural radiance fields used to render 3D scenes. DINo learns an additional latent variable $\bm{\alpha}_{t}$ and corresponding neural ODE $\dot{\bm{\alpha}}={\bm{f}}_{\psi}(\bm{\alpha})$, such that the true system is decoded using ${\bm{g}}_{\bm{\theta}}(\bm{\alpha}_{t},{\bm{x}})$, where ${\bm{g}}_{\bm{\theta}}$ is the INR. Encoding is performed using ‘auto-decoding‘, where $\bm{\alpha}$ is optimized with gradient-based methods to invert ${\bm{g}}_{\bm{\theta}}$. The benefit of this approach is that, like PINNs, we do not need to discretize the spatial domain. However, unlike PINNs, the LVM formulation is more expressive and $\bm{\alpha}$ many different systems to be learned, for example different initial boundary conditions. A similar architecture is the continuous reduced-order modeling approach (CROM, Chen et al. (2023a)), which also combines neural ODEs and INRs, but unlike DINo uses a learned discretization-dependent encoder during learning, rather than auto-decoding.
More recently, significant progress has been made in machine learning for weather prediction using deep latent variable models. While these models are primarily data-driven and do not have physically-interpretable latent spaces, their performance has been achieved using several small but crucial physics-informed design decisions that we will discuss here.
@@
Weather forecasting traditionally relies on computationally intensive physically-derived models, often combined in ensembles, to mitigate uncertainties arising from chaotic atmospheric dynamics and measurement limitations. Despite advancements, the accuracy of these methods for medium-range, regional, and global climate modeling is dictated by the computational resources and model fidelity. Additionally, they cannot leverage extensive historical weather data beyond optimizing model parameters with system identification. Moreover, the complexity of weather dynamics requires combining models for each component (e.g. clouds, pollution, etc), and jointly tuning these models to improve net prediction accuracy is challenging/ Machine learning approaches offer promise by directly predicting measurements from historical datasets, circumventing some computational challenges. Recent years have witnessed a significant advancement in machine learning methods for weather prediction, enabled by large capacity models, hardware accelerators, and clean historical datasets (Voosen, 2023). We discuss the design of three prominent nonlinear latent variable models in this context. An empirical analysis of deep learning-based weather prediction can be found in Karlbauer et al. (2024).
@@
Bi et al. (2023) proposed Pangu-Weather. Like FourCastNet, this model builds off vision transformers, but instead of using AFNOs, proposes a novel three-dimensional (3D) Earth-specific transformer architecture (3DEST) to better model weather correlation with latitude, longitude, and altitude. The key design decision is in the positional encoding used by the 3DEST, which encodes absolute position coordinates rather than relative ones and incorporates translation invariance. This encoding uses 527 times more embedding parameters, but the authors did not observe issues with prediction or optimization. The model also employs ‘hierarchical temporal aggregation’ to address error accumulation in long-horizon forecasts. To mitigate this source of error, the authors learn models that are trained on 1-, 3-, 6- and 24-hour intervals so that long-horizon forecasts can be computed using an aggregation of models, using 3-4 predictions at most. However, with each 3D ViT having 64M parameters, this aggregation strategy results in a 256M parameter model, which presents engineering issues at deployment. Similarly, Schmude et al. (2024) proposed Prithvi-WxC, a ViT-based approach with multi-axis attention to facilitate long-range interaction with limited attention window sizes and, to mitigate block boundary artifacts, an attention window shifting method was introduced. The model uses 160 input variables in order to address a more diverse set of downstream tasks. To reason over this many inputs, it uses both local and global masking and attention mechanisms during training to learn to model local and global phenomena. However, the resulting model has 2.3B parameters, posing significant challenges to deployment.
@@
Lam et al. (2023) introduced GraphCast, a graph neural network (GNN) approach for medium-range forecasting. GraphCast builds on the work of Pfaff et al. (2021) learning mesh-based simulations using graph neural networks. GraphCast utilizes a hierarchical graph structure over the globe, resembling 3DEST, with six levels of resolution. The model’s encoder projects Earth-oriented grid measurements onto this multi-mesh representation while the decoder aggregates nearby mesh values back to the grid. Operating on a 40,962-node multi-mesh, the GNN conducts simulation via simultaneous message passing at various spatial resolutions. Despite making 6-hour predictions, during training, it is evaluated autoregressively for up to 3 days. Remarkably, GraphCast can predict a 10-day forecast in under a minute at 0.25 global resolution with approximately 40 million parameters. Comparative evaluations suggest GraphCast outperforms Pangu-Weather across most measurement modalities, possibly due to the GNN’s inductive bias, which resembles finite-element PDE solvers and propagates computations coherently in space. This behavior contrasts with ViT architecture, which needs latent self-attention mechanisms to learn such computations. However, the graph architecture does complicate some aspects of forecasting, such as downscaling and regional prediction.
Further extensions of this work use diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) for decoding. GenCast (Price et al., 2025) extends GraphCast by using a graph transformer for the latent dynamics and a conditional diffusion model for decoding. Compared to GraphCast, which has a deterministic decoder, sampling GenCast produces diverse high-fidelity forecasts. GraphCast is less capable of such high-fidelity predictions since the deterministic decoder requires averaging over any present stochasticity. Precipitation diffusion (PreDiff, Gao et al. (2023)) also uses a conditional diffusion model for decoding weather predictions. Using a VAE to encode several steps of observations, the conditional diffusion model is trained to forecast a short lookahead window. An interesting aspect of this model is incorporating physical plausibility into predictions. For image generation, classifer-based guidance (Dhariwal and Nichol, 2021) uses the gradient of a pretrained classifier to improve the quality of the image generation. Similarly, Gao et al. (2023) replaces the classifier with knowledge-guidance, for example, a function that scores samples that reflect energy conservation, and uses this function to guide samples towards these physically plausible characteristics. The authors use energy conservation guidance in an $n$-body simulation experiment and anticipated precipitation intensity for a precipitation forecast experiment.
@@
Despite progress, several challenges persist in current weather forecasting models. A significant hurdle is their reliance on ‘reanalysis’ data derived from complex Bayesian state estimation processes, which combine meteorological measurements with simulation models. This reliance on processed data prevents these ML models from forecasting in real time using raw data as desired. Additionally, challenges include scaling models to go beyond the 0.25 training data and match the high spatial resolution of 0.1 of top NWP systems, extending forecasting horizons, and increasing model capacity. There is also the question of the extent to which physics-based elements from NWP can be combined with ML models to enhance their performance, given that NWP forecasts are inherently physically plausible, while current ML-based forecasts lack this feature. This question is an active topic of study, e.g., (Kochkov et al., 2024). The next section explores examples of incorporating physics-informed elements into model learning.
3.6 Bridging Domain Knowledge and Function Approximation for System Identification
@@
System identification approaches lie on a spectrum depending on modeling assumptions. If a significant amount of domain knowledge is assumed, the dynamics model will be highly structured. If correct, it will generalize well but will most likely be inflexible to unanticipated observed phenomena. On the other hand, one could model with no domain knowledge and use a black-box function approximator [^3] One could argue that using many popular function approximators, such as Gaussian processes, come with a smoothness assumption. These models have no generalization guarantees and, therefore, require expansive datasets and regularization to ensure good modeling accuracy across the domain of interest. To achieve an effective combination of these two approaches, one would need to balance structural biases that facilitate generalization with adaptivity to the data that prevents the biases from leading to underfitting (Ljung, 2010; Lutter et al., 2021). This section reviews these approaches, which build on models discussed earlier.
@@
Physics-informed neural networks can incorporate an additional dataset as an additional loss term in the objective. In addition to residual and boundary loss terms, the combined objective (equation (19)) has an additional dataset loss similar to the boundary loss term (Raissi et al., 2017a; b). Moreover, since the PINNs objective is quite general to any differentiable function approximator, Wang et al. (2021d) extend DeepONets by adding the PINNs residual loss to the data-driver operator loss as a way of encoding physics knowledge.
@@
3.7 Physics-inspired Model Invariances
Physical models in the form of ODEs, PDEs, or algebraic equations are often invariant to specific types of operations or transformations. For instance, many dynamical processes such as molecular, rigid body, elasto or fluid dynamics, chemical reaction kinetics, etc. are modeled under the assumption of energy conservation and the laws of thermodynamics (Cueto and Chinesta, 2023). This also translates to continuum material models, which should be invariant to the orientation of the coordinate system and consider the orientation of the material’s microstructure. Likewise, in computer vision, object detection should be invariant to the position, orientation, and scale of the object. Graph-based representations, e.g., in the context of molecular interactions or particle systems (Sanchez-Gonzalez et al., 2020), should often also be invariant to rotations, translations, reflections, and permutations. To discover invariants and conservation quantities from datasets using machine learning, Wetzel et al. (2020) proposed a “Siamese twin” neural network architecture, which aims to classify data points as related to the same invariant value. Similarly, Ha and Jeong (2021) introduced an intuitive noise-variance loss function for training a NN to classify data points related and non-related to the same invariant value.
Mathematically defining and incorporating relevant invariance and equivariances into machine learning models has led to several advancements in applying machine learning to complex physical domains, e.g., (Cohen and Welling, 2016; Fuchs et al., 2020; Satorras et al., 2021; Hoogeboom et al., 2022; Weiler et al., 2023; Thiemann et al., 2025). Denoting a set of transforms as $\mathsf{T}$, the $\mathsf{T}$-invariance of a function or operator can be expressed as:
@@ 3.7 Physics-inspired Model Invariances
(model-based) $f_{\bm{\theta}}({\bm{x}})$ $=f_{\bm{\theta}}(T({\bm{x}}))\quad\forall T\in\mathsf{T},$ (42)
(feature-based) $f_{\bm{\theta}}(\phi({\bm{x}}))$ $=f_{\bm{\theta}}(\phi(T({\bm{x}})))\quad\forall T\in\mathsf{T},$ (43)
@@
In summary, material modeling is subject to several types of invariance requirements such as thermodynamic consistency, rotational invariances, and functional requirements. It has already been demonstrated through various techniques that this prior knowledge can be successfully incorporated into ML frameworks for constitutive modeling. In particular, ML models can be fully physics-augmented, i.e., the requirements can be embedded into the model architectures by careful choice of features, outputs, and NN structure. This additional structure generally greatly improves the model performance, generalization, robustness, and interpretability and reduces data requirements.
@@
Symmetries and energy conservation are also important aspects of the modeling of dynamic systems. As already discussed in section 3.2, the equations of motion of a dynamical system can be formulated in an energy-conserving, i.e., thermodynamically consistent, fashion by deriving them through the Lagrangian or Hamiltonian formalisms (Lutter and Peters, 2023; Cranmer et al., 2020b; Greydanus et al., 2019). In these approaches, the invariance is model-based and incorporated through the model conception as the (Lagrangian ${\mathcal{L}}$, Hamiltonian ${\mathcal{H}}$, …) potentials are learned/parameterized by a NN – and not the dynamics themselves in the form of an ODE, e.g., through a neural ODE, see section 3.1. This concept has also been applied to Koopman operators (section 3.5), for the data-based discovery of conservation laws in Kaiser et al. (2018) and could be further extended to neural operators. Likewise, enforcing conservation laws can also be achieved in the context of neural solution fields (Richter-Powell et al., 2022), c.f. section 3.3, neural operators (Liu et al., 2023b; 2024b), c.f. section 3.4, or neural processes (Hansen et al., 2023), c.f. section 4.3.
However, energy conservation is usually an idealization of real system behavior, and dissipative effects are present in most real-world systems. Dissipation behavior can be considered in a similar fashion through application of formalisms such as the Port-Hamiltonian (Desai et al., 2021; Neary and Topcu, 2023; Roth et al., 2025), see equation (16), pseudo-Hamiltonian (Eidnes et al., 2023), or GENERIC (General Equation for the Non-Equilibrium Reversible–Irreversible Coupling) frameworks (González et al., 2019; Hernández et al., 2021), which is structurally very similar to the GSM framework mentioned above for material modeling.
@@ 4.1 Multi-task Learning
Thanasutives et al. (2021) propose an MTL approach of solving the original PDE along with an auxiliary related PDE with different coefficients. In this work, they integrate MTL at the architecture and optimization levels. For the architecture, cross-stitch (Misra et al., 2016) is employed to share information between two parallel baseline networks for original and auxiliary tasks. At each layer, a cross-stitch unit is used to linearly combine the input activations from the two networks to allow flexible sharing of representations between the two branches. The authors evaluate two MTL techniques for learning, an uncertainty-weighting loss (Kendall et al., 2017) and gradient surgery (Yu et al., 2020). The uncertainty-weighted physics-informed loss function (Uncert) is defined as a combination of task-specific PINNs loss functions weighted by the uncertainty of each PDE solution. The loss function is defined as,
@@
where $N$ refers to the number of tasks, $\sigma_{i}$ is a gradient-based trainable parameter that indicates the uncertainty of each PDE solution, and $\mathcal{L}_{\textrm{PINN}}$ is a similar loss function as in equation (19). “Gradient surgery” is implemented using PCGrad (Yu et al., 2020) to modify the average of the unweighted PINNs loss functions for mitigating conflicting gradients between tasks. PCGrad projects the loss gradient of task $i$ onto the normal vector of task $j$’s loss gradient and vice versa. The model parameters are updated using $\delta_{\textrm{PC}}\theta$ shown below,
@@
where ${\bm{g}}^{i}_{\textrm{PC}}$ is the modified gradient of task $i$ after applying PCGrad. The proposed algorithm provides a better generalization on highly nonlinear domains in terms of lower error rates on unseen data points in comparison to related baselines.
Zou and Karniadakis (2023) propose a new PINN architecture that suites MTL. In this work, a multi-head architecture is presented, named multi-head physics-informed neural network (MH-PINN), where a body $\Phi$ is shared by a collection of tasks (e.g., ODEs/ PDEs) while task-specific heads $H_{k},k=1,...,N$ are learned separately for each model. The body consists of nonlinear hidden layers, while linear layers are used for the different heads. Consequently, the body provides a set of basis functions for each head to generate task-specific solutions. The body $\Phi:\mathbb{R}^{d_{{\bm{x}}}}\rightarrow\mathbb{R}^{d}$ is a function parameterized with parameter $\theta$, and each head $H_{k}$ is a linear function. The task-specific solution is defined as $\hat{{\bm{u}}}_{k}({\bm{x}})=H_{k}\circ\Phi(x),\forall x\in\Omega$. Given an MTL dataset $\{\mathcal{D}_{k}\}_{k=1}^{N}$, the MTL loss function is defined as follows:
@@
Similarly, Wang et al. (2023) ease the training of PINNs by employing a multi-task optimization paradigm where multiple auxiliary tasks are solved together with a main task. Knowledge transfer from one task facilitates solving others. The problem of automatic traffic modeling is considered, which has multiple modalities such as traffic speed and density. Given several datasets measuring different traffic scenarios and MTL over datasets and modalities, a similar PINN architecture, as in Zou and Karniadakis (2023), is trained with a shared body network and only a final layer per task. This architecture design significantly increases the data available to train the initial PINN layers, which build an internal feature representation. The MTL ensures that this feature representation is useful across traffic conditions and modalities. As a drawback, the model requires an adaptive training scheme in order for the multi-task aspect to be effective, which significantly increases the implementation complexity.